OpenStudy (anonymous):

The fourth roots of -8+8i√3 using DeMoivre's Theorem

OpenStudy (shubhamsrg):

take 16 common out and try to see if it helps

ganeshie8 (ganeshie8):

that helps u in writing the number in polar from, start wid that @darkserious

OpenStudy (anonymous):

polar form as in z = a + bi?

ganeshie8 (ganeshie8):

\[-8 + 8\sqrt{3}i = 16 e^{i\frac{2\pi}{6}}\]

ganeshie8 (ganeshie8):

4th roots are : \[ \left( 16 e^{i\frac{2\pi}{6} + 2k\pi}\right)^{\frac{1}{4}}\] \(k=0,1,2,3\)